ANALYSIS OF THE ROTATIONAL STRUCTURE IN THE HIGH-RESOLUTION INFRARED SPECTRA OF cis, cis- AND trans, trans-1,4-DIFLUOROBUTADIENE-1-d_{1} AND trans, trans-1,4-DIFLUOROBUTADIENE-1,4-d_{2}

\underline{NORMAN C. CRAIG}, YIHUI CHEN, YUHUA LU, CHRISTOPHER F. NEESE, and DEACON J. NEMCHICK, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; THOMAS A. BLAKE, Pacific Northwest National Laboratory, Richland, WA 99352.

Samples of cis,cis- and trans,trans-1,4-difluorobutadiene-1-d_{1} (DFBD) and trans,trans-DFBD-1,4-d_{2} have been synthesized and investigated with high-resolution (0.0015 cm^{−1}) infrared spectroscopy. For the first two species the rotational structure in more than one band has been analyzed. For the 1,4-d_{2} species the spectrum of only one C-type band was available in an isotopic mixture. Ground state rotational constants are reported for all three molecules. It is proposed that quartic centrifugal distortion constants computed with a B3LYP/cc-pVTZ model can be used to assess the quality of observed rotational constants. The favorable comparison of predicted and observed ground state rotational constants for all four 13C species of cis,trans-DFBD, which is MW active,^{a} demonstrates that the ground state rotational constants for the ^{13}C species of the cis,cis and trans,trans isomers can be successfully predicted with high accuracy. Rotational constants for a full set of isotopologues will be used to determine accurate semiexperimental equilibrium structures of the cis,cis and trans,trans species of DFBD.

aN. C. Craig, C. M. Oertel, D. C. Oertel, M. J. Tubergen, R. J. Lavrich, A. M Chaka J. Phys. Chem. A 106, 4230-4235 (2002).